OpenStudy (anonymous):

derivative of (arc sec(u))/(3u + 2)

OpenStudy (anonymous):

You should post your question as "I'm confused and need help with this problem" rather than just putting the question, we take the code of conduct seriously.

OpenStudy (anonymous):

ok sorry but whats the purpose?

OpenStudy (anonymous):

d/du arcsec(u) = 1/sqrt(1-u^2) d/du 1/(3u+2) = d/du (3u+2)^(-1) = (-1)(3u+2)^(-2) (3) = -3/(3u+2)^2 d/du [ arcsec(u) / (3u+2)] = [d/du arcsec(u)] (1/(3u+2)] + arcsec (u) [d/du (1/(3u+2)] = 1/sqrt(1-u^2)(3u+2) - 3arcsec(u) /(3u+2)^2

OpenStudy (anonymous):

Sometimes people try and cheat on here, and openstudy is about studying together, not getting answers. (: Anyway follow that for the rest of the problems similar to this.

OpenStudy (anonymous):

ok thanks